OpenStudy (anonymous):

Teresa has a copy of her school directory on her desk. Students whose last names start with the letters A-F occupy the first 5 pages of the book. Those students with last names G-M occupy the next 7 pages, those with last names N-R the following 10 pages, and those with last names S-Z the final 10 pages. If she opens up the directory and picks out a student at random, what is the probability that this student’s last name begins with a letter from the first half of the alphabet?

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

I think its 7/32

OpenStudy (anonymous):

it's 12/32

OpenStudy (anonymous):

because the first half of the alphabet ends at the letter "N"

OpenStudy (anonymous):

is their a smaller fraction to put that

OpenStudy (texaschic101):

I am not good at these...but lets see what I can do. There are a total of (5 + 7 + 10 + 10) = 32 pages Half of the alphabet end at letter M. So we have 5 pages + 7 pages.....till halfway through the alphabet. Leaving us with : 12/32 = 3/8

OpenStudy (anonymous):

okay thanks guys

OpenStudy (texaschic101):

anytime :)